40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 3 23:09:34 2017, awade, DailyProgress, optic, Moved FSS EOM backward 6 cm 100_0289.JPG100_0290.JPG100_0294.JPG
    Reply  Fri Jan 6 13:38:04 2017, rana, HowTo, FSS, Debugging the FSS loops 
       Reply  Tue Jan 10 12:04:30 2017, awade, HowTo, FSS, Debugging the FSS loops 7x
          Reply  Wed Jan 11 09:01:37 2017, rana, HowTo, FSS, Debugging the FSS loops 
       Reply  Fri Feb 3 11:48:54 2017, awade, HowTo, FSS, Debugging the FSS loops SPandRMSNorthandSouthErrSig.pdfSPandRMSNorthandSouthErrSig_copy.pdfdata20170203_SpectraErrorandFastMonSignals.zip
          Reply  Sun Feb 26 20:42:07 2017, awade, HowTo, FSS, Debugging the FSS loops 
             Reply  Mon Feb 27 20:03:16 2017, awade, HowTo, FSS, Debugging the FSS loops 20170227_SPNorthPDRF_D980454.zipD980454_NorthRFPD_PSDNoise_wrong.pdf
                Reply  Thu Mar 2 13:56:25 2017, awade, HowTo, FSS, Debugging the FSS loops batchplot20170228_SpecNorthPD_DCandAC.pdf20170228_SPNorthPDRF_D980454_RedoneWithCorrectPort.zip
Message ID: 1799     Entry time: Tue Jan 3 23:09:34 2017     Reply to this: 1804
Author: awade 
Type: DailyProgress 
Category: optic 
Subject: Moved FSS EOM backward 6 cm 

I was unable to remove a ~32 kHz ringing from the north path lock. I thought possibly that it might have to do with polarisation alignment into the EOM used to actuate in the FSS lock: this in the past had converted phase acutation into amplitude acutation from misalignment. However, the BB-EOM was placed so close to the prior waveplate that it was imposible to get a PBS in to check.  I've moved the BB-EOM 6 cm backward to give space for future diagnostics. Picture before and after attached. It seemed to be aligned ok.

After some realignment down the path I was able to get a lock again,  but there is still ringing, this time at 22 kHz.  The waveform and frequency change with some choices of gains. 

I've attached screen shots of what I seen on the osciloscope, not that that helps much. Not sure if a transfer function would help, if the loop is ringing that doesn't make the measurment easy. This is probably something obvious but I'm leaving it for today.

---

FYI side note, there is now an obvious small spray of light that I can see on the CCD cameras, like a flare.  Tuning up the alignment seemed to move it off a bit, I don't remember seeing it there before.  There doesn't seem to be an obvious scattering before the vacuum can.

Attachment 1: 100_0289.JPG  474 kB  | Hide | Hide all
100_0289.JPG
Attachment 2: 100_0290.JPG  533 kB  | Hide | Hide all
100_0290.JPG
Attachment 3: 100_0294.JPG  496 kB  | Hide | Hide all
100_0294.JPG
ELOG V3.1.3-